# Private Heterogeneous Federated Learning Without a Trusted Server Revisited: Error-Optimal and Communication-Efficient Algorithms for Convex Losses

### v2

@@ 3.2 Error-Optimal and Computationally Efficient ISRL-DP FL for Non-Smooth Losses
Theorem 3.4 (Non-smooth FL via subgradient).
@@
The above subgradient complexity improves over the state-of-the-art gradient complexity (Lowy and Razaviyayn, 2023b) for an ISRL-DP FL algorithm with optimal excess risk. Lowy and Razaviyayn (2023b) apply Nesterov smoothing to ISRL-DP MB-SGD. As discussed earlier, implementing the smoothing approach is computationally costly. Moreover, the results in Lowy and Razaviyayn (2023b) assume i.i.d. silo data. The subgradient complexity result in Theorem 3.4 also improves over the gradient complexity of Algorithm 1 with smoothing.
An alternative approach is to use the convolutional smoothing technique (Kulkarni et al., 2021) to optimize non-smooth functions. For certain regimes, this approach improves the gradient complexity over the subgradient method, with a worse communication complexity compared to the Nesterov smoothing approach. See Appendix D for details.
Depending on the FL application, communication or computation may be more of a bottleneck. If communication efficiency is more of a priority, the smoothed version of Algorithm 1 should be used. On the other hand, if computational efficiency is more pressing, Algorithm 4 or Algorithm 6 is recommended.
@@ 4 Numerical Experiments
We validate our theoretical findings with numerical experiments on MNIST data. As shown in Figures 3 and 3, our algorithm consistently outperforms Lowy and Razaviyayn (2023b). We use a similar experimental setup to Lowy and Razaviyayn (2023b), as outlined below.
@@
We compare our algorithm against the One-pass ISRL-DP MB-SGD of Lowy and Razaviyayn (2023b). Recall that Lowy and Razaviyayn (2023b) did not provide theoretical guarantees for their multi-pass ISRL-DP MB-SGD with heterogeneous silos.
@@ 5 Concluding Remarks and Open Questions
We have studied private federated learning in the absence of a trusted server. We characterized the minimax optimal excess risk bounds for heterogeneous ISRL-DP FL, answering an open question posed by Lowy and Razaviyayn (2023b). Further, our algorithms advanced the state-of-the-art in terms of communication and computational efficiency. For smooth losses, the communication complexity of our optimal algorithm matches the non-private lower bound.
@@ Appendix A Further Discussion of Related Work
There is a large and growing body of work on DP optimization. Most of this work focuses on the centralized setting, with Lipschitz convex loss functions in $\ell_{2}$ geometry [Bassily et al., 2014, 2019, Feldman et al., 2020, Zhang et al., 2022] and $\ell_{p}$ geometry [Asi et al., 2021, Bassily et al., 2021]. Recently, we have started to learn more about other central DP optimization settings, such as DP optimization with non-uniformly Lipschitz loss functions/heavy-tailed data [Lowy and Razaviyayn, 2023a], non-convex loss functions [Gao and Wright, 2023, Lowy et al., 2024], and min-max games [Boob and Guzmán, 2023]. There has also been work on the interactions between DP and other ethical desiderata, like fairness [Lowy et al., 2023b] and robustness [Wu et al., 2023], as well as DP optimization with side access to public data [Lowy et al., 2023c]. Despite this progress, much less is known about DP distributed optimization/federated learning, particularly in the absence of a trustworthy server.
@@
There have been many works attempting to ensure privacy of people’s data during the federated learning (FL) process. Some of these works have utilized user-level differential privacy [McMahan et al., 2018, Geyer et al., 2017, Levy et al., 2021], which can be practical for cross-device FL with a trusted server. Several works have also considered inter-silo record-level DP (ISRL-DP) or similar notions to ensure privacy without a trusted server [Heikkilä et al., 2020, Liu et al., 2022, Lowy and Razaviyayn, 2023b, b, Zhou and Chowdhury, 2023].
The state-of-the-art theoretical bounds for convex ISRL-DP FL are due to Lowy and Razaviyayn [2023b] which gave minimax error-optimal algorithms and lower bounds for the i.i.d. setting, and suboptimal algorithms for the heterogeneous setting. We close this gap by providing optimal algorithms for the heterogeneous setting. Additionally, we improve over the communication complexity and gradient complexity bounds in Lowy and Razaviyayn [2023b].
@@ Appendix B Multi-Stage Implementation of ISRL-DP Accelerated MB-SGD
Algorithm 5 Multi-stage Accelerated Noisy MB-SGD [Lowy and Razaviyayn, 2023b]
@@
We will need the following result for the excess risk bound, which is due to [Lowy and Razaviyayn, 2023b] (Theorem F.1).
@@ C.1 Gradient Complexity of Algorithm 1 under Different Parameter Regimes
Appendix D Optimize Non-Smooth Losses via Convolutional Smoothing
@@ D.2 Algorithm and Analysis
Theorem D.5 (Non-smooth FL via convolutional smoothing).
@@
We present the complete proof below. Let $\beta:=L\sqrt{d}/r$. By properties of the smoother (D.2), we know that $\tilde{f}_{s}$ is convex, $L$-Lipschitz and $\beta$-smooth. We can thus reuse Lemma C.3, restated below.
@@
To prove this, it suffices to show that the results Lemma B.1 hold for the multi-stage implementation of Algorithm 7. The proof is essentially the same as that in Lowy and Razaviyayn [2023b] by observing that the variance of the estimate is still $O(L^{2}/m)$ by Theorem D.4.
@@
Plugging $\beta$ into Equation 24, we obtain the gradient complexity of $\widetilde{O}\left(M^{3/2}n^{3/4}+M^{5/4}n^{11/8}\right)$.
@@
Our algorithm for non-smooth functions uses Lemma E.1 and Algorithm 1 as follows: First, property (1) above allows us to apply Algorithm 1 to optimize $f_{\beta}$ and obtain an excess population risk bound for $f_{\beta}$, via Theorem 2.1. Inside Algorithm 1, we will use property (3) to compute the gradient of $f_{\beta}$. Then, property (2) enables us to extend the excess risk guarantee to the true function $f$, for a proper choice of $\beta$.
@@
$K_{i}\geq\max\left(1,\frac{\varepsilon n}{4\sqrt{2R\ln(2/\delta)}}\right)$, $R_{i}=\min\left(Mn_{i},\frac{M\varepsilon^{2}n_{i}^{2}}{d}\right)+1$, and $\sigma_{i}^{2}=\frac{256L^{2}R_{i}\ln(\frac{2.5R}{\delta})\ln(2/\delta)}{n_{i}^{2}\varepsilon^{2}}$ for $i\in[k]$in Algorithm 4. Then, Algorithm 4 is $(\varepsilon,\delta)$-ISRL-DP and achieves the following excess risk bound:
@@
Privacy. The privacy analysis of Lowy and Razaviyayn [2023b, Theorem D.1] for ISRL-DP MB-SGD holds verbatim in the nonsmooth case when we replace subgradients by gradients: the only property of $f(\cdot,x)$ that is used in the proof is Lipschitz continuity. Thus, Algorithm 3 is $(\varepsilon,\delta)$-ISRL-DP if the noise variance is $\sigma^{2}\geq\frac{256L^{2}R\ln(2.5R/\delta)\ln(2/\delta)}{\varepsilon^{2}n^{2}}$. By our choice of $\sigma_{i}^{2}$, we see that phase $i$ of Algorithm 4 is $(\varepsilon,\delta)$-ISRL-DP on data $\{B_{i,l}\}_{l=1}^{N}$. Since the batches $\{B_{i,l}\}_{i=1}^{\tau}$ are disjoint for all $l\in[N]$, the full Algorithm 4 is $(\varepsilon,\delta)$-ISRL-DP by parallel composition [McSherry, 2009].
@@
Gradient complexity. Recall $K_{i}\geq\max\left(1,\frac{\varepsilon n_{i}}{4\sqrt{2R_{i}\ln(2/\delta)}}\right)$. Choosing the minimum $K_{i}$, we have

### v3

@@ 3.2 A Direct Subgradient Algorithm
Theorem 3.5 (Nonsmooth FL via subgradient).
@@
The above subgradient complexity improves over the previous state-of-the-art gradient complexity (Lowy et al., 2023a) for an ISRL-DP FL algorithm with optimal excess risk. Lowy et al. (2023a) apply Nesterov smoothing to ISRL-DP MB-SGD. As discussed earlier, implementing the smoothing approach is computationally costly. Moreover, the results in (Lowy et al., 2023a) assume i.i.d. silo data.
The precise statement and proof of Theorem 3.5 can be found in Theorem G.1.
@@ 4 Numerical Experiments
We validate our theoretical findings with numerical experiments on MNIST data. As shown in Figures 3 and 3, our algorithm consistently outperforms Lowy et al. (2023a). We use a similar experimental setup to Lowy et al. (2023a), as outlined below.
@@
We compare our algorithm against the One-pass ISRL-DP MB-SGD of Lowy et al. (2023a). Recall that Lowy et al. (2023a) did not provide theoretical guarantees for their multi-pass ISRL-DP MB-SGD with heterogeneous silos.
@@ 5 Concluding Remarks and Open Questions
We have studied private federated learning in the absence of a trusted server. We characterized the minimax optimal excess risk bounds for heterogeneous ISRL-DP FL, answering an open question posed by Lowy et al. (2023a). Further, our algorithms advanced the state-of-the-art in terms of communication and computational efficiency. For smooth losses, the communication complexity of our optimal algorithm matches the non-private lower bound.
@@ Appendix A Further Discussion of Related Work
There is a large and growing body of work on DP optimization. Most of this work focuses on the centralized setting, with Lipschitz convex loss functions in $\ell_{2}$ geometry [Bassily et al., 2014, 2019, Feldman et al., 2020, Zhang et al., 2022] and $\ell_{p}$ geometry [Asi et al., 2021, Bassily et al., 2021]. Recently, we have started to learn more about other central DP optimization settings, such as DP optimization with non-uniformly Lipschitz loss functions/heavy-tailed data [Lowy and Razaviyayn, 2023b], non-convex loss functions [Gao and Wright, 2023, Lowy et al., 2024], and min-max games [Boob and Guzmán, 2023]. There has also been work on the interactions between DP and other ethical desiderata, like fairness [Lowy et al., 2023b] and robustness [Wu et al., 2023], as well as DP optimization with side access to public data [Lowy et al., 2023c]. Despite this progress, much less is known about DP distributed optimization/federated learning, particularly in the absence of a trustworthy server.
@@
There have been many works attempting to ensure privacy of people’s data during the federated learning (FL) process. Some of these works have utilized user-level differential privacy [McMahan et al., 2018, Geyer et al., 2017, Levy et al., 2021], which can be practical for cross-device FL with a trusted server. Several works have also considered inter-silo record-level DP (ISRL-DP) or similar notions to ensure privacy without a trusted server [Heikkilä et al., 2020, Liu et al., 2022, Lowy et al., 2023a, a, Zhou and Chowdhury, 2023].
The state-of-the-art theoretical bounds for convex ISRL-DP FL are due to Lowy et al. [2023a] which gave minimax error-optimal algorithms and lower bounds for the i.i.d. setting, and suboptimal algorithms for the heterogeneous setting. We close this gap by providing optimal algorithms for the heterogeneous setting. Additionally, we improve over the communication complexity and gradient complexity bounds in Lowy et al. [2023a].
@@ Appendix B Multi-Stage Implementation of ISRL-DP Accelerated MB-SGD
Algorithm 5 Multi-stage Accelerated Noisy MB-SGD [Lowy et al., 2023a]
@@
We will need the following result for the excess risk bound, which is due to [Lowy et al., 2023a] (Theorem F.1).
@@ C.1 Gradient Complexity of Algorithm 1 under Different Parameter Regimes
Appendix D Optimize Nonsmooth Losses via Convolutional Smoothing
@@ D.2 Algorithm and Analysis
Theorem D.5 (Nonsmooth FL via convolutional smoothing).
@@
We present the complete proof below. By properties of the smoother (D.2), we know that $\tilde{f}_{s}$ is convex, $L$-Lipschitz and $\beta$-smooth. We can thus reuse Lemma C.3, restated below.
@@
To prove this, it suffices to show that the results Lemma B.1 hold for the multi-stage implementation of Algorithm 7. The proof is essentially the same as that in Lowy et al. [2023a] by observing that the variance of the estimate is still $O(L^{2}/m)$ by Theorem D.4.
@@
Plugging $\beta$ into Equation 25, we obtain the gradient complexity (in expectation) of $\widetilde{O}\left(M^{3/2}n^{3/4}+M^{5/4}n^{11/8}\right)$.
@@
Our algorithm for nonsmooth functions uses Lemma E.1 and Algorithm 1 as follows: First, property (1) above allows us to apply Algorithm 1 to optimize $f_{\beta}$ and obtain an excess population risk bound for $f_{\beta}$, via Theorem 2.1. Inside Algorithm 1, we will use property (3) to compute the gradient of $f_{\beta}$. Then, property (2) enables us to extend the excess risk guarantee to the true function $f$, for a proper choice of $\beta$.
@@
$K_{i}\geq\max\left(1,\frac{\varepsilon n_{i}}{4\sqrt{2R_{i}\ln(2/\delta)}}\right)$, $R_{i}=\min\left(Mn_{i},\frac{M\varepsilon^{2}n_{i}^{2}}{d}\right)+1$, and $\sigma_{i}^{2}=\frac{256L^{2}R_{i}\ln(\frac{2.5R}{\delta})\ln(2/\delta)}{n_{i}^{2}\varepsilon^{2}}$ for $i\in[k]$in Algorithm 4. Then, Algorithm 4 is $(\varepsilon,\delta)$-ISRL-DP and achieves the following excess risk bound:
@@
Privacy. The privacy analysis of Lowy et al. [2023a, Theorem D.1] for ISRL-DP MB-SGD holds verbatim in the nonsmooth case when we replace subgradients by gradients: the only property of $f(\cdot,x)$ that is used in the proof is Lipschitz continuity. Thus, Algorithm 3 is $(\varepsilon,\delta)$-ISRL-DP if the noise variance is $\sigma^{2}\geq\frac{256L^{2}R\ln(2.5R/\delta)\ln(2/\delta)}{\varepsilon^{2}n^{2}}$. By our choice of $\sigma_{i}^{2}$, we see that phase $i$ of Algorithm 4 is $(\varepsilon,\delta)$-ISRL-DP on data $\{B_{i,l}\}_{l=1}^{N}$. Since the batches $\{B_{i,l}\}_{i=1}^{\tau}$ are disjoint for all $l\in[N]$, the full Algorithm 4 is $(\varepsilon,\delta)$-ISRL-DP by parallel composition [McSherry, 2009].
@@
Subgradient complexity. Recall $K_{i}\geq\max\left(1,\frac{\varepsilon n_{i}}{4\sqrt{2R_{i}\ln(2/\delta)}}\right)$. Choosing the minimum $K_{i}$, we have